OpenStudy (anonymous):

If a piece of ribbon is 5 and 1 over 2. feet long, how long is it in inches? 33 inches 55 inches 60 inches 66 inches

OpenStudy (anonymous):

will fan and medal

OpenStudy (igreen):

\(5 \dfrac{1}{2}\) feet long?

OpenStudy (anonymous):

yes

OpenStudy (igreen):

Okay, well 1 foot = 12 inches. So: \(5 \dfrac{1}{2} \cdot 12 =~?\)

OpenStudy (anonymous):

what does that mean

OpenStudy (igreen):

The dot is multiplication, can you multiply 5.5 * 12?

OpenStudy (anonymous):

ok give me a sec i will

OpenStudy (anonymous):

66

OpenStudy (igreen):

Yep, so that's your answer.

OpenStudy (anonymous):

thx alot can u help me a couple more plx